The sum of the ages of two brothers is 38 years 4 years ago the age of the Elder brother was the square of the younger Brothers age. Find their ages.
Answer by Alan3354(69443)

The sum of the ages of two brothers is 38 years. 4 years ago the age of the Elder brother was the square of the younger Brothers age. Find their ages.
----------------
e + y = 38 now
4 years ago, e + y = 30
-----------
e + y = 30
e = y^2
---> 5 & 25 then
9 & 29 now
===============
The hard way:
e + y = 30
e = y^2
y^2 + y - 30 = 0
(y+6)*(y-5) = 0
y = 5
etc